# AGE-US: automated gestational age estimation based on fetal ultrasound images

### v1

@@ Abstract
Keywords:
Fetal Ultrasound Segmentation Deep Learning Distance Map
@@ 1 Introduction
This work proposes AGE-US, a method for automated GA estimation based on US images of the head, abdomen, and femur planes acquired in the third trimester of pregnancy. Our method uses transfer learning and distance maps for structures where segmentation masks are scarce, making it suitable for smaller image datasets. The main contributions of this study include (1) a single-encoder dual-decoder U-Net architecture that reduces the complexity of the segmentation model without statistically significant performance degradation, and (2) a method for addressing the unavailability of femur segmentation masks by incorporating distance maps. Finally, to provide a more interpretable alternative to end-to-end GA estimation methods [7], AGE-US integrates these contributions in a fully automated pipeline that mimics standard clinical practice.
@@ Acknowledgments
This work was supported by the Galician Ministry of Culture, Education, Professional Training and University (ED481A-2023-121, 2024-2027 ED431G2023/04, 2022-2025 ED431C2022/19, and GRC2021/48). The previous grants are co-funded by the European Regional Development Fund (ERDF/FEDER program). This work is also supported by the Spanish Ministry of Science and Innovation through the project COLLAGE PID2023-149549NB-I00; and by the Spanish Ministry of Science and Innovation, the Spanish Research Agency, and the European Union through the “Ramón y Cajal” 2022 program (RYC2022-035469-I and RYC2022-035960-I).

### v2

@@ 1 Introduction
This work proposes AGE-US [^1] Source code and model weights: https://gitlab.citius.gal/cesar.diaz.parga/age-us, a method for automated GA estimation based on US images of the head, abdomen, and femur planes acquired in the third trimester of pregnancy. Our method uses transfer learning and distance maps for structures where segmentation masks are scarce, making it suitable for smaller image datasets. The main contributions of this study include (1) a single-encoder dual-decoder U-Net architecture that reduces the complexity of the segmentation model without statistically significant performance degradation, and (2) a method for addressing the unavailability of femur segmentation masks by incorporating distance maps. Finally, to provide a more interpretable alternative to end-to-end GA estimation methods [7], AGE-US integrates these contributions in a fully automated pipeline that mimics standard clinical practice.
@@ Acknowledgments
This work was supported by the Galician Ministry of Culture, Education, Professional Training and University (ED481A-2023-121, 2024-2027 ED431G2023/04, 2022-2025 ED431C2022/19, and GRC2021/48). The previous grants are co-funded by the European Regional Development Fund (ERDF/FEDER program). This work is also supported by the Spanish Ministry of Science and Innovation through the project COLLAGE PID2023-149959OA-I00; and by the Spanish Ministry of Science and Innovation, the Spanish Research Agency, and the European Union through the “Ramón y Cajal” 2022 program (RYC2022-035469-I and RYC2022-035960-I).